Shift report 20030516 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri May 16 2003 - 23:55:45 EDT

  • Next message: Apache: "Shift report 20030517 08:00-16:00"
    ------------------------------------------------------------
    Date: Fri May 16 23:55:45 2003, Shift: 16:00-24:00
    Supervisor: JHL, with: 
    Log book#14 pages 111 - 111
    
    Shift Summary:
    ------------------------------------------------------------
    Beam: 
    No beam for the entire shift.
    A few attemps after recovering from today's access  
    didn't make a store due to poor polarization.  
    And variety of problems:
    Linac problems, Quench Link Interlock, Corrector Magnet Supply Problems...
    
    Plan:
    MRS 90deg. 1050A 20K trigger3
    MRS 90deg. 350A 50K trigger3
    MRS 90deg. 0Field 15K trigger3
    ...
    FS 4deg. 1/8A 40K trigger2
    FS 4deg. 1/1A 105K trigger2
    
    
    Detector Problems:
    Si has been turned off (preamp power) due to some problems (unknown?).
    
    To prevent a potential problem:
    Magnets should be in "stanby" mode while beam is off 
    for an extensive time to prevent over-heating.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri May 16 2003 - 23:56:12 EDT